Grade 12Physical Chemistry

An ideal solution was obtained by mixing methanol and ethanol .if the partial vapour pressure of methanol and ethanl are 2.619 kPa and 4.556 kPa reapec.the coposition of vapour (in termas of mole fration ) will be.?

Shriya Mehrotra
11 Years agoGrade 12

2 Answers

Ramreddy
11 Years ago
Calculate total pressure from the given individual pressures.
P total= 2.619 + 4.556
= 6.725 kPa
p methanol= (total pressure) * Xmethanol
Xmethanol=
Xethanol can be calculated from in similar way.
........................................................................................
….….….….….….….
Momina batool
9 Years ago
First thing to keep in mindP(any component)=total P×mole fraction of that component.Total p=(2.619+4.556)kPa =7.175Now using the above given relation2.619=7.175×nn~.365So mole fraction of other component will be 1-.365=.635Best of luck with the question☺️